A couple is planning to have three children.What is the probability that?


1
Expert's answer
2021-02-28T16:41:33-0500

There are 8 possible arrangements of girls and boys. For example, GGB means the first two children are girls and the third child is a boy. All 8 arrangements are (approximately) equally alike.

The possible arrangements are:

GGG

GGB

GBG

GBB

BGG

BGB

BBG

BBB

Assume that the probability for aither B or G is "\frac{1}{2}"

So, the probability of any of these arrangements is:

"\frac{1}{2} \times \frac{1}{2} \times \frac{1}{2} = \frac{1}{8}"